I am wanting to check an answer.
I am suppose to use the rational zero test to find all the rational zeros of f(x).
f(X)=2x^3+3x^2-11x-6
I got -2, -4, -3 is this correct?

Answer by vasumathi(46)

Solution : f(X)=2x^3+3x^2-11x-6
Plug in x = 2 in this
f(2) = 2(2)^3 + 3(2)^2-11(2)-6 = 16+12-22-6=28-28 = 0
so x = 2 is a root
divide the polynomial with (x-2)
Then we get
2x^2+7x+3
Plug in x = -3 in this
2(-3)^2 +7(-3)+3
=18-21+3
=0
so x = -3 is another root
and on dividing 2x^2+7x+3 with x+3 we get 2x+1
and the root of 2x+1 = 0 is x = -0.5
so the roots are x = 2,-3,-0.5